HOW DIFFERENT ARE ORTHO AND PARA HYDROGEN? A COMPUTATIONAL EXAMINATION OF POTENTIALS, BOUND STATES, AND SCATTERING OF HYDROGEN COMPLEXES

\underline{KELLY J. HIGGINS}, Spectral Sciences, Inc., Burlington, MA 01803; WILLIAM KLEMPERER, Department of Chemistry and Chemical Biology, Harvard University, Cambridge, MA 02138.

The interaction between molecular hydrogen and other molecules is studied through ab initio calculations with a focus on differences in behavior between ortho and para hydrogen. These differences manifest themselves in various ways, including in the rotational spectra of complexes^{a}^{,}^{b} and in pressure broadening cross sections.^{c} Two specific systems, H_{2}-HCN and H_{2}-OCS, will be examined. For H_{2}-HCN, a new 4-D potential surface will be presented along with bound state and scattering calculations, while for H_{2}-OCS new calculations utilizing a previously presented potential^{d} will be discussed. These two systems display markedly different ground state geometries that lead to different behavior of the ortho and para hydrogen complexes. In addition to the specific systems a more general discussion will examine the effect of potential anisotropy and ground state geometry on the properties of hydrogen complexes and how the differences between ortho and para interactions can be used.

aZ. Yu, et al., J. Chem. Phys. 127, 054305 (2007).

bM. Ishiguro, et al., J. Chem. Phys. 115, 5155 (2001).

cM. Mengel, D.C. Flatin, and F. C. De Lucia, J. Chem. Phys. 112, 4069 (2000).

dK.J. Higgins, Z. Yu, and W. Klemperer, 62nd International Symposium on Molecular Spectroscopy (2007).